The results of the nationwide aptitude test in mathematics are normally distributed with mean=63 and standard deviation =5 . What is the 97.5th percentile?

1
Expert's answer
2022-02-28T17:17:32-0500

"P(Z<z)=0.975\to z=1.96."

"\frac{x-63}{5}=1.96\to x=1.96*5+63="72.8.